OpenStudy (anonymous):

The graph of the function f(x)=(k-1)x^2+(k+1)x+k-1 touches the x-axis at only one point.If k>1,find a)the value of k. @rational

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

if this answer is correct then i'll tell you how i did it. either k = 1/3 or k=3

OpenStudy (rational):

that means f(x) has only one zero, so the discriminant has to be \(0\) : \[b^2-4ac=0\]

OpenStudy (rational):

\[f(x)=(k-1)\color{blue}{x}^2+(k+1)\color{blue}{x}+k-1\] \(a=k-1\) \(b=k+1\) \(c=k-1\) plugin the values and solve \(k\)

OpenStudy (anonymous):

Answer will be k=3. Thnx @rational :)